OpenStudy (anonymous):

A physicist found that a force of 0.72 N was measured between two charged spheres. The distance between the spheres was 0.90 m. Calculate the force between the two spheres when the distance was decreased to 0.50 m. F2 = _____.

OpenStudy (anonymous):

coulomb's law states that \[F \alpha \frac{ Q_1Q_2 }{ d^2 }\] F=Force between the charges Q(1)= quantity of charge on the first particle Q(2)=quantity of charge on the second particle d=distance apart